# Towards Federated Graph Learning in One-shot Communication

### v5

@@ 4.3 Global Pseudo-Graph Generation
$N^{c}=\sum_{k}^{m}N_{k}^{c},\quad\boldsymbol{\mu}^{c}=\frac{1}{N^{c}}\sum_{k}^{m}N_{k}^{c}\boldsymbol{\mu}_{k}^{c},$ (9)
$\boldsymbol{s^{2}}^{c}=$ $\frac{1}{N^{c}-m}(\sum_{k}^{m}(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c}+\sum_{k}^{m}N_{k}^{c}(\boldsymbol{\mu}_{k}^{c}-\boldsymbol{\mu}^{c})^{2}),$
@@
Figure 5: Analysis of a client on the CiteSeer dataset. Left: The performance difference of our method (ours) and two variants (ft and ft_dist) compared with $M_{G}$ trained solely on the global pseudo-graph. Middle: The value of $H$ and $w_{dist}$ of each class. Right: The value of $H$ is influenced by both the number of samples and their homophily.
Inspired by the graph condensation (Liu et al., 2022; Xiao et al., 2024), we generate a small-size global pseudo-graph $G^{\prime}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ by aligning its feature distribution with the aggregated global distribution. This pseudo-graph serves as a representative of the overall graph data across all clients. The generation process is illustrated in the top of Figure 3. We firstly pre-set the number of nodes of each class in the pseudo-graph and initialize the trainable node features $\mathbf{X}^{\prime}$ from Gaussian noise. The adjacency matrix $\mathbf{A}^{\prime}$ is defined using a trainable link predictor $g_{\theta}$:
@@
Following (Xiao et al., 2024), to optimize $\mathbf{X}^{\prime}$ and $g_{\theta}$, we first propagate node features $\mathbf{X}^{\prime}$ as in Eq. 5. Then we calculate the class-wise sample mean $\boldsymbol{\mu}^{\prime c}$ and sample variance $\boldsymbol{s^{2}}^{\prime c}$ of the propagated node features $\mathbf{X}^{\prime prop}$ following Eq. 6. Then we calculate the alignment loss:
@@
We could optimize the $\mathbf{X}^{\prime}$ and $g_{\theta}$ to generate the global pseudo-graph by the overall optimization objective:
$\min_{\mathbf{X}^{\prime},\theta}\ (L_{align}+\alpha L_{smt}).$ (13)
Once the global pseudo-graph is generated, the server distributes it to clients for further personalized training.
Privacy-preservation with Secure Aggregation. We demonstrate compatibility with Secure Aggregation protocols in uploading and aggregating $\{\boldsymbol{\mu}_{k}^{c},\boldsymbol{s^{2}}_{k}^{c},N_{k}^{c}\}$. Secure Aggregation enables the server to compute the sum of large, user-held data vectors securely, without accessing individual client contributions. This property aligns well with the aggregation process in Eq. 9, as it can be expressed as a series of weighted average operations. The calculation of $N^{c}$ and $\boldsymbol{\mu}^{c}$ are inherently weighted averages. The calculation of $\boldsymbol{s^{2}}^{c}$, while more complex, can also be decomposed into weighted average operations:
$\boldsymbol{s^{2}}^{c}$ $=\frac{\sum_{k}^{m}(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c}}{N^{c}-m}$ (14)
$+\frac{\sum_{k}^{m}N_{k}^{c}(\boldsymbol{\mu}_{k}^{c})^{2}-2\boldsymbol{\mu}^{c}\sum_{k}^{m}\boldsymbol{\mu}_{k}^{c}+N^{c}\sum_{k}^{m}(\boldsymbol{\mu}^{c})^{2}}{N^{c}-m}.$
@@
Table 1: Performance on datasets under Louvain partition with 10 clients. OOM represents the out-of-memory issue.
@@
4.4 Personalized Training with Node Adaptive Distillation
With the downloaded global pseudo-graph, each client trains its personalized model locally. However, the imbalanced local graph data often results in biased models that perform well on major classes but neglect minor classes. To achieve better personalization (high accuracy) and generalization (high F1-macro), we propose a two-stage adaptive personalized training approach with node-adaptive distillation to leverage both the global information in the global pseudo-graph and personal information in the local data.
Stage 1
Each client trains a local model $M_{G}$ using only the global pseudo-graph, ensuring $M_{G}$ to capture global, balanced knowledge. The training objective is to minimize:
A copy of the resulting $M_{G}$ is detached as a fixed teacher model for the next stage. $L_{ce}$ denotes the cross-entropy loss.
Stage 2
$L_{2^{nd}}=L_{ft}+L_{dist},$ (16)
$\gamma_{i}$ is a node-specific weighting factor that controls the balance between local fine-tuning and global knowledge distillation. To determine $\gamma_{i}$ for each node, we utilize the soft labels $\tilde{\mathbf{y}}_{i}$ obtained in Section 4.2 and $\mathbf{w_{dist}}$ in Eq. 3:
$\gamma_{i}=\beta\tilde{\mathbf{y}}_{i}\cdot\mathbf{w_{dist}},$ (18)
where $\beta$ is a scaling hyper-parameter. Intuitively explained, if a node is predicted to class $c$ with a small $w_{dist}(c)$ (e.g. node $b$ in Figure 4), it likely belongs to major classes with high homophily. In this case, $\gamma_{i}$ is small, and $M_{i}$ primarily learns through supervised fine-tuning; If a node is predicted to class $c$ with a large $w_{dist}(c)$ or has a blended prediction (e.g. node $a$ in Figure 4), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus $\gamma_{i}$ should be large to incorporate corresponding global knowledge from $M_{G}$.
@@ 5.1 Experimental Setup
We conduct our experiments on 12 datasets considering both transductive and inductive learning settings. For the transductive setting, we use Cora, CiteSeer, and PubMed from (Yang et al., 2016) and ogbn-arxiv and ogbn-products (Hu et al., 2020). We also use Computers, Photo, CS, Physics datasets (Shchur et al., 2018). For the inductive setting, we use Flickr (Zeng et al., 2019), Reddit (Hamilton et al., 2017), and Reddit2 (Zeng et al., 2019). A summary of these datasets is provided in Table 4 in the Appendix. To simulate distributed subgraphs, we employ two advanced graph partitions: Louvain-based Label Imbalance Split and Metis-based Label Imbalance Split (Li et al., 2024c).
@@
We compare our method with 8 state-of-the-art methods including 4 pFGL methods [^1] We do not include FedSage (Zhang et al., 2021) and AdaFGL (Li et al., 2024a) since they need extra communication rounds to train neighborhood generators or federated optimization. They cannot train models within one-shot communication.: (1) FedPUB (Baek et al., 2023), (2) FedGTA (Li et al., 2024b), (3) FedTAD (Zhu et al., 2024b), (4) FedSpray (Fu et al., 2024), which pays attention to minority, and 2 OFL methods: (5) DENSE (Zhang et al., 2022), (6) Co-Boost (Dai et al., 2024). We also include (7) Standalone, which locally trains models, and (8) FedAvg with fine-tuning, as additional baselines.
Metrics.
We employ both accuracy(%) and F1-macro(%) metrics. F1-macro averages F1 score across all classes, providing a more robust measurement of the model’s generalization, particularly under imbalanced data. we report the mean and standard deviation across three runs.
Hyper-parameters and Configurations.
We limit the communication round to a single round. We employ the two-layer GCN (Kipf & Welling, 2016) with the hidden dimension set to 64. Detailed settings and configurations under model heterogeneity are described in the Appendix.
@@ 5.2.1 Performance Comparison
The performance under the Louvain partition with 10 clients is shown in Table 1 and results with more clients and under the Metis partition are in the Appendix. Our method consistently outperforms on both accuracy and F1-macro, demonstrating the superiority of our method in deriving models with better personalization and local generalization. FedTAD encounters the OOM issue due to its graph diffusion operation. FedSpray focuses on minor classes, but it requires numerous communication for optimization, offering little advantage in one-shot communication. Furthermore, our method even generally outperforms other methods with 100 communication rounds, which is shown in Section A.2.2 in Appendix. More results on Computers, Photo, CS, and Physics datasets are also listed in Appendix.
@@ 5.2.2 Ablation Study
We conduct the ablation study to analysis our method. We compare our method with 5 variants: (1) ft: only fine-tuning $M_{G}$ to obtain $M_{i}$ on local graph data, (2) ft_focal: fine-tuning with focal loss (Ross & Dollár, 2017) to handle data imbalance, (3) ft_dist: in the second stage, performing distillation on all nodes with a fixed $\gamma_{i}$; (4) nd_ft_dist: performing node adaptive distillation along with fine-tuning, without HRE, (5) ours: performing the complete method (including HRE), and (6) hre_nd_dist: performing our complete method but without the first stage in personalized training, instead train a model from scratch in the second stage.
The results under the Louvain partition and Metis partition are shown in Figure 6 and Figure 8 in the Appendix, respectively. The focal loss cannot effectively address the non-IID graph data since it’s hard to determine the proper weights. Fixed distillation cannot balance improvements in accuracy and F1 macro as it does not distinguish the majority from the minority. For nodes in major classes, the global knowledge of $M_{G}$ may have a negative impact, as demonstrated by the left sub-figure in Figure 5. This motivates us to develop node-adaptive distillation. The node adaptive distillation is more flexible to determine when to introduce the global knowledge of $M_{G}$ in fine-tuning. And the HRE strategy further improves performance by generating a better global pseudo-graph with more precise estimations. Comparing ours and hre_nd_dist, we conclude that $M_{G}$ provides a necessary start-point for fine-tuning. Furthermore, We justify the rationale for considering homophily in determining the major classes by $H$ values in Eq. 2 in the right panel of Figure 5. The full analysis is in the Appendix.
@@
Louvain Cora CiteSeer ogbn-arxiv
@@
Figure 7: Communication costs versus performance.
5.2.3 Performance under Model Heterogeneity
We conduct experiments on 10 clients with heterogeneous models. Under model heterogeneity, most pFGL methods do not work since they rely on the model parameter aggregation. The results under Louvain and Metis partitions are presented in Table 3 and Table 11 in the Appendix, respectively. The results demonstrate the superior performance of our method in both accuracy and F1-macro.
5.2.4 Inductive Performance
The results on inductive graph datasets under the Louvain partition are shown in Table 2 and more results are shown in the Appendix. Since Flickr has lower homophily, we use 2-layer SIGN (Frasca et al., 2020) for better graph learning. Our method is scalable and generally performs best.
@@ 5.2.5 Communication Analysis
We analyze the communication costs both theoretically and empirically. Our method involves uploading the class-wise statistics ($O(Chf)$) and downloading the generated global pseudo-graph ($O(n^{\prime}f+n^{\prime 2})$, where $n^{\prime}$ is the number of nodes in the global pseudo-graph). Both communications are independent of the model size. Figure 7 reports the communication costs of uploading and downloading on a client on small-scale Cora and large-scale ogbn-products graph datasets under the Metis partition. Note that we report the performance and costs of 1, 10, and 100 communication rounds for FedAvg, FedPUB, FedGTA, and FedSpray. The results demonstrate that our method achieves better performance with lower communication costs. Detailed analysis of communication, computation, and memory is shown in Table 19, Table 20, and Table 21 in the Appendix.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
A.1 Experimental Setup
@@
Dataset Description
We conduct experiments on 12 graph datasets and the summary is listed in Table 4. Our selected datasets cover a wide range from 1) small-scale to large-scale, 2) transductive to inductive, and 3) academic domain to industry domain.
Additional Baselines Description
In DENSE and Co-Boost, the generators are trained to generate node features, and the topology structure is constructed using the $K$-Nearest Neighbors strategy as outlined in (Zhu et al., 2024b). Note that we perform additional fine-tuning processes on FedTAD, DENSE, and Co-Boost to ensure better performance.
Hyper-parameters and Model Configurations
In our one-shot personalized federated graph learning setting, we limit the communication round to a single round. For each client, we employ a two-layer GCN (Kipf & Welling, 2016) with the hidden layer dimension set to 64 by default. For baseline methods, the number of local training epochs is tuned from 3 to 100, and hyper-parameters are set as recommended in their respective papers or fine-tuned for optimal overall performance. For our method, we keep the size of the generated global pseudo-graph as small as possible. On small-scale Cora, CiteSeer and PubMed datasets, we set the number of nodes in each class in the global pseudo-graph as 1. We set $p$ as 0.25%, 0.04%, 1%, 1%, 1%, 0.5%, 0.2%, 0.05%, 0.05% on ogbn-arixv, ogbn-products, Computers, Photo, CS, Physics, Flickr, Reddit, and Reddit2 datasets. We tune $f_{th}$ within [0.95, 1] and tune $d_{th}$ according to the scale of the graphs. We tune the $\beta$ within [0.1, 1] to control the range of $\gamma_{i}$. For other hyper-parameters in global pseudo-graph generation, we adopt the setting in (Xiao et al., 2024). Under model heterogeneity, we adopt 10 different GNN models categorized into two types: coupled GNNs and decoupled GNNs. For coupled GNNs, we adopt 6 models which are: 2-layer GCN with 64, 128, 256, and 512 hidden dimension, 3-layer GCN with 64 and 512 hidden dimension. For decoupled GNNs, we adopt APPNP (Gasteiger et al., 2018), SSGC (Zhu & Koniusz, 2021), SGC ($K$=2) and SGC ($K$=4) (Wu et al., 2019).
Experimental Environments
The experimental machine is with Intel(R) Xeon(R) CPU E5-2680 v4@2.40GHz, and 6$\times$×NVIDIA GeForce RTX 3090. The operating system is Ubuntu 20.04 and the version of PyTorch is 2.1.2.
@@
A.2 More Experimental Results and Full Analysis
A.2.1 Performance Comparison
We first conduct experiments in the transductive learning setting and report the accuracy and F1-macro metrics. The datasets range from small-scale graphs to large-scale graphs with millions of nodes. We also varies the number of clients to evaluate the stability of methods. With 10 clients, the experimental results under Louvain and Metis partitions are shown in Table 1 and Table 6, respectively. With 20 clients, the experimental results under Louvain and Metis partitions are shown in Table 5 and Table 8, respectively. Our method consistently outperforms others in terms of accuracy and F1-macro. Under the non-IID scenarios, pFGL methods show marginal or even negative performance gain over Standalone, highlighting the ineffectiveness of current pFGL methods in personalizing models on non-IID graph data in a single communication round. The state-of-the-art OFL methods (DENSE and Co-Boost) also have difficulty generating high-quality pseudo-data by the ensemble model when each model is trained on non-IID local graph data. FedTAD encounters OOM when handling million-scale graphs due to its graph diffusion operation, while our method remains scalable. FedGTA performs worst because its aggregation is based on the unreliable local models’ predictions under non-IID data in a single round. FedSpray focuses on minor classes, but it requires numerous communication for optimization, offering little advantage over others in one-shot communication.

### v6

@@ 4.3 Global Pseudo-Graph Generation
$N^{c}=\sum_{k=1}^{m}N_{k}^{c},\quad\boldsymbol{\mu}^{c}=\frac{1}{N^{c}}\sum_{k=1}^{m}N_{k}^{c}\boldsymbol{\mu}_{k}^{c},\quad\boldsymbol{s^{2}}^{c}=\frac{1}{N^{c}-m}(\sum_{k=1}^{m}(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c}+\sum_{k=1}^{m}N_{k}^{c}(\boldsymbol{\mu}_{k}^{c}-\boldsymbol{\mu}^{c})^{2}),$ (9)
@@
Inspired by the graph condensation [35, 51], we generate a small-size global pseudo-graph $G^{\prime}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ by aligning its feature distribution with the aggregated global distribution. This pseudo-graph serves as a representative of the overall graph data across all clients. The generation process is illustrated in the top of Figure 2. We first pre-set the number of nodes of each class in the pseudo-graph and initialize the trainable node features $\mathbf{X}^{\prime}$ from Gaussian noise. The adjacency matrix $\mathbf{A}^{\prime}$ is derived using a trainable link predictor $g_{\theta}$:
@@
Following [51], to optimize $\mathbf{X}^{\prime}$ and $g_{\theta}$, we first propagate node features $\mathbf{X}^{\prime}$ as in Eq. 5. Then we calculate the class-wise sample mean $\boldsymbol{\mu}^{\prime c}$ and sample variance $\boldsymbol{s^{2}}^{\prime c}$ of the propagated node features $\mathbf{X}^{\prime prop}$ following Eq. 6. Then we calculate the alignment loss:
@@
We could optimize $\mathbf{X}^{\prime}$ and $g_{\theta}$ to generate the global pseudo-graph by the overall optimization objective: $\min_{\mathbf{X}^{\prime},\theta}\ (L_{align}+\alpha L_{smt})$. Once the optimization finishes, the server distributes $G^{\prime}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ to clients for further personalized training.
Privacy-preservation with Secure Aggregation. We demonstrate compatibility with Secure Aggregation protocols in uploading and aggregating $\{\boldsymbol{\mu}_{k}^{c},\boldsymbol{s^{2}}_{k}^{c},N_{k}^{c}\}$. Secure Aggregation is a crucial technique in federated learning that enables the server to compute the sum of large, user-held data vectors securely, without accessing individual client contributions. This property aligns well with the aggregation process in Eq. 9, as it can be expressed as a series of weighted average operations. The calculations of $N^{c}$ and $\boldsymbol{\mu}^{c}$ are inherently weighted average operations. The calculation of $\boldsymbol{s^{2}}^{c}$, while more complex, can also be decomposed into weighted average operations:
$\boldsymbol{s^{2}}^{c}=\frac{\sum_{k}^{m}(N_{k}^{c}-1)\boldsymbol{s^{2}}_{k}^{c}}{N^{c}-m}+\frac{\sum_{k}^{m}N_{k}^{c}(\boldsymbol{\mu}_{k}^{c})^{2}-2\boldsymbol{\mu}^{c}\sum_{k}^{m}\boldsymbol{\mu}_{k}^{c}+N^{c}\sum_{k}^{m}(\boldsymbol{\mu}^{c})^{2}}{N^{c}-m}.$ (13)
@@
4.4 Personalized Training with Node Adaptive Distillation
With the downloaded global pseudo-graph, each client trains its personalized model locally. However, the imbalanced local graph data often results in biased models that perform well on major classes but neglect minor classes. To achieve better personalization ($\uparrow$) and generalization (F1$\uparrow$), we propose a two-stage personalized training approach with node-adaptive distillation to leverage both global information in the global pseudo-graph and personal information in the local graph.
Stage 1
Each client trains a generalized model $M_{G}$ using only the global pseudo-graph, ensuring $M_{G}$ to capture global, balanced knowledge. The training objective is to minimize:
A copy of $M_{G}$ is detached as a fixed teacher model for the next stage. $L_{ce}$ denotes the cross-entropy loss. Note that if the server knows the model architecture, an alternative is to conduct this stage on the server. However, we can adapt to more rigorous scenarios where models’ intellectual property needs to be protected. In those cases, it’s necessary to conduct this stage on each client.
Stage 2
$L_{2^{nd}}=L_{ft}+L_{dist},$ (15)
where $\gamma_{i}$ is a node-specific weighting factor to control the balance between local fine-tuning and global knowledge distillation. To determine $\gamma_{i}$ for each node, we re-use the soft labels $\tilde{\mathbf{y}}_{i}$ obtained in Section 4.2 and $\mathbf{w}_{dist}$ in Eq. 3:
$\gamma_{i}=\beta\tilde{\mathbf{y}}_{i}\cdot\mathbf{w}_{dist},$ (17)
where $\beta$ is a scaling hyper-parameter. Intuitively explained, if a node is predicted to class $c$ with a small $w_{dist}(c)$ (e.g. node $b$ in Figure 7), it likely belongs to major classes with high homophily. In this case, $\gamma_{i}$ is small, and $M_{i}$ primarily learns through supervised fine-tuning; If a node is predicted to class $c$ with a large $w_{dist}(c)$ or has a blended prediction (e.g. node $a$ in Figure 7), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus $\gamma_{i}$ should be large to incorporate the corresponding global knowledge from $M_{G}$.
Table 1: Performance on datasets under Louvain partition with 10 clients.
@@ 5.1 Experimental Setup
We conduct experiments on 12 datasets, including both transductive learning datasets: Cora, CiteSeer, PubMed [56], ogbn-arxiv, ogbn-products [23], Computers, Photo, CS, Physics datasets [40], and inductive learning datasets: Flickr [59], Reddit [19], and Reddit2 [59]. Table 4 summarizes these datasets. We employ two advanced graph partitions to simulate distributed subgraphs: Louvain-based Label Imbalance Split and Metis-based Label Imbalance Split [34].
@@
We compare our method with 9 state-of-the-art methods, including 5 pFGL methods [^1] We do not include FedSage [63] and AdaFGL [32] since they need extra communication rounds to train neighborhood generators or federated optimization. They cannot train models within one-shot communication.: (1) FedPUB [2], (2) FedGTA [33], (3) FedTAD [69], (4) FedSpray [15], (5) FedLoG [30], and 2 OFL methods: (6) DENSE [62], (7) Co-Boost [9]. We also include (8) Standalone, which locally trains models, and (9) FedAvg with fine-tuning. Detailed descriptions are in the Appendix A.7.
Configurations.
We limit the communication round to a single round. We employ the two-layer GCN [31] with the hidden dimension set to 64. We employ both accuracy(%) and F1-macro(%) metrics. F1-macro averages the F1 scores across all classes, providing a more robust measurement of the model’s generalization, particularly under imbalanced data. We report the mean and standard deviation across three runs. Detailed configurations are described in the Appendix A.2.
@@
Figure 4: Ablation study under Metis partition.
Figure 5: Client-level analysis on CiteSeer dataset. Left: Comparison of different personalized training methods. $M_{G}$ is the generalized model trained in Stage 1. Middle: The value of $H$ and $w_{dist}$ for each class. Right: The value of $H$ is determined by both samples quantity and homophily.
@@ 5.2.1 Performance Comparison
The performance under the Louvain partition with 10 clients is shown in Table 1, and results with more clients and results under the Metis partition are shown in the Appendix A.3.1. Our method consistently outperforms on both accuracy and F1-macro, demonstrating the superiority of our method in deriving models with better personalization and local generalization. FedTAD encounters the OOM issue due to its graph diffusion operation, and FedLoG encounters the OOM issue due to its $k$-hop subgraph construction for each node. FedSpray and FedLoG also focus on minor and underrepresented classes, but they require numerous communications for optimization under non-IID scenarios, offering little advantage in one-shot communication. Furthermore, our method (1 communication round) even outperforms other multi-round federated learning methods (100 communication rounds) on average, which is shown in the Appendix A.3.2. More results on Computers, Photo, CS, and Physics datasets are listed in the Appendix A.3.6. We also include the experiments with other prevalent graph models (GraphSage [19], GAT [45], and SGC [49]) in Appendix A.3.7.
@@ 5.2.2 Ablation Study
We conduct the ablation study to analyze our method. We compare our method with 4 variants: (1) ft: only fine-tuning $M_{G}$ to obtain $M_{i}$ on local graph data, (2) ft(focal): fine-tuning with focal loss [39] to handle data imbalance, (3) ft(fixed $\gamma$): in the second stage, performing distillation with a fixed $\gamma_{i}$ for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE, and (5) ours: performing the complete method (including HRE).
The results under the Louvain partition and Metis partition are shown in Figure 4 and Figure 4, respectively. The focal loss cannot effectively address the non-IID graph data since it’s hard to determine the proper weights. Fixed $\gamma$ cannot balance improvements in accuracy and F1-macro as it cannot distinguish the majority from the minority. For nodes in major classes, the global knowledge of $M_{G}$ may have a negative impact, as demonstrated by the left sub-figure in Figure 5. This motivates us to develop node-adaptive distillation, which is more flexible to determine when to introduce global knowledge of $M_{G}$ in fine-tuning. HRE strategy further improves performance by generating a better global pseudo-graph with more precise estimations. Furthermore, we justify the rationale for considering homophily in determining the major classes by $H$ values in Eq. 2 in the right panel of Figure 5. The detailed analysis is in the Appendix A.3.3.
@@
Method Cora CiteSeer ogbn-arxiv
@@
Figure 6: Communication costs.
5.2.3 More Experiments
Inductive Performance.
The results on inductive datasets under the Louvain partition are shown in Table 2. More results are shown in the Appendix A.3.5. Our method is scalable and outperforms.
Model Heterogeneity.
We conduct experiments on 10 clients with heterogeneous models and configurations are described in Appendix A.2. Under model heterogeneity, most pFGL methods do not work since they rely on the model parameter aggregation. The results under Louvain and Metis partitions are presented in Table 3 and Table 11, respectively. The results demonstrate the superior performance of our method in both accuracy and F1-macro. Details are in the Appendix A.3.4.
Hyper-parameters Analysis.
We study our method’s sensitivity to the hyper-parameters in Appendix A.3.8. Our method is generally robust under a proper range of hyper-parameters values.
@@ 5.2.4 Communication Analysis
We analyze the communication costs both theoretically and empirically. Our method involves uploading the class-wise statistics ($O(Chf)$) and downloading the generated global pseudo-graph ($O(n^{\prime}f+n^{\prime 2})$, $n^{\prime}$ is the pre-set number of nodes in the global pseudo-graph, which is small as described in Appendix A.2). Both are independent of the model size. Figure 6 reports the communication costs of baseline methods and our method on the Cora datasets under the Louvain and Metis partition. The results show that our method has lower communication costs. We include the detailed theoretical analysis of communication, computation, and memory complexity in the Appendix A.4. We also introduce an equivalent server-efficient variant of our method in the Appendix A.5.
@@ Appendix A Appendix
A.1 Illustration for Determining $\gamma_{i}$ in Node Adaptive Distillation
We illustrate how to determine $\gamma_{i}$ for each node in each client’s local graph. We pre-compute the vector $\mathbf{w}_{dist}\in\mathbb{R}^{C}$ by Eq. 3 and soft labels by Label Propagation in Section 4.2. Considering soft label of each node, if a node is predicted to class $c$ with a small $w_{dist}(c)$ (e.g. node $b$ in Figure 7), it likely belongs to major classes with high homophily. In this case, $\gamma_{i}$ is small, and $M_{i}$ primarily learns through supervised fine-tuning; If a node is predicted to class $c$ with a large $w_{dist}(c)$ or has a blended prediction (e.g. node $a$ in Figure 7), it likely belongs to minor classes or has low homophily. These nodes are often neglected during supervised fine-tuning. Thus $\gamma_{i}$ should be large to incorporate the corresponding global knowledge from $M_{G}$.
A.2 Experimental Setup
@@
Dataset Description.
We conduct experiments on 12 graph datasets, and the summary is listed in Table 4. Our selected datasets cover a wide range from 1) small-scale to large-scale, 2) transductive to inductive, and 3) academic domain to industry domain.
Hyper-parameters and Model Configurations.
In our one-shot personalized federated graph learning setting, we limit the communication round of the 5 pFGL methods and the FedAvg method to a single round. For 2 OFL methods (DENSE and Co-Boost), the generators are trained to generate node features, and the topology structure is constructed using the $K$-Nearest Neighbors strategy as outlined in [69]. Note that we perform additional fine-tuning processes on FedAvg, FedTAD, DENSE, and Co-Boost to ensure better performance.
For each client, we employ a two-layer GCN [31] with the hidden layer dimension set to 64 by default. For baseline methods, the number of local training epochs is tuned from 1 to 100, and hyper-parameters are set as recommended in their respective papers or fine-tuned for optimal overall performance. For our method, we keep the size of the generated global pseudo-graph as small as possible. On small-scale Cora, CiteSeer, and PubMed datasets, we set the number of nodes in each class in the global pseudo-graph to 1. For other datasets, we set the number of nodes in each class to $p$ percent of the total number of nodes of the corresponding class in the training set. We set $p$ as 0.25%, 0.04%, 1%, 1%, 1%, 0.5%, 0.2%, 0.05%, 0.05% on ogbn-arxiv, ogbn-products, Computers, Photo, CS, Physics, Flickr, Reddit, and Reddit2 datasets. We use the Adam optimizer and tune the learning rate. We tune $f_{th}$ within [0.95, 1] and tune $d_{th}$ according to the scale of the graphs. We tune the $\beta$ within [0.1, 1] to control the range of $\gamma_{i}$. For other hyper-parameters in global pseudo-graph generation, we adopt the setting in [51].
Under model heterogeneity, we adopt 10 different GNN models categorized into two types: coupled GNNs and decoupled GNNs. For coupled GNNs, we adopt 6 models, which are: 2-layer GCN with 64, 128, 256, and 512 hidden dimensions, and 3-layer GCN with 64 and 512 hidden dimensions. For decoupled GNNs, we adopt APPNP [16], SSGC [68], SGC ($K$=2) and SGC ($K$=4) [49].
For metrics, we employ both accuracy(%) and F1-macro(%) metrics to comprehensively evaluate the performance. F1-macro averages the F1 scores across all classes, providing a more robust measurement of the model’s generalization, particularly under imbalanced data. For each experiment, we report the mean and standard deviation across three runs.
Experimental Environments.
The experimental machine is with Intel(R) Xeon(R) CPU E5-2680 v4@2.40GHz, and 6$\times$×NVIDIA GeForce RTX 3090. The operating system is Ubuntu 20.04, and the version of PyTorch is 2.1.2. We use PyG 2.5.2 [11].
@@
A.3 More Experimental Results and Detailed Analysis
A.3.1 Performance Comparison
We first conduct experiments in the transductive learning setting and report the accuracy and F1-macro metrics. The datasets range from small-scale graphs to large-scale graphs with millions of nodes. We also vary the number of clients to evaluate the stability of methods. With 10 clients, the experimental results under Louvain and Metis partitions are shown in Table 1 and Table 6, respectively. With 20 clients, the experimental results under Louvain and Metis partitions are shown in Table 5 and Table 8, respectively. Our method consistently outperforms others in terms of accuracy and F1-macro. Under the non-IID scenarios, pFGL methods show marginal or even negative performance gain over Standalone, highlighting the ineffectiveness of current pFGL methods in personalizing models on non-IID graph data in a single communication round. The state-of-the-art OFL methods (DENSE and Co-Boost) also have difficulty generating high-quality pseudo-data by the ensemble model when each model is trained on non-IID local graph data. FedTAD encounters OOM when handling million-scale graphs due to its graph diffusion operation. FedLoG encounters OOM issues due to its node-wise $k$-hop subgraph construction. In contrast, our method remains scalable. FedGTA performs badly because its aggregation is based on the unreliable local models’ predictions under non-IID data in a single round. FedSpray focuses on minor classes, but it requires numerous communications for optimization, offering little advantage over others in one-shot communication. FedLoG focuses on local generalization, but it cannot work well under severe non-IID scenarios within one-shot communication.